5. Use Laplace transforms to solve the following value problem. (10pts) (a) y - y = e cost y(0)=1 yo =-1
Answer #1

5. La Here the De y-y = e cost y (0) = 1, Y (o)=-1 a & {y } & lest yit) at 22 [est was to - Jesset yay at - Tē©ycey - 1093-(S-ilt +(S-1) S6 (smut Sint OSO Sint at T2 -soigt cost -(S-1)t costat e Since el 2 (0-0) - (3-1) pē = (s-41-8 +1] - (-131 1Now , ང་ ། ( ད ༤༥༡) (༩༤ - ) བྱ S ༡ s(༩s- ད༥) (s) བན ད -+ : sf - ད 2 sl cong 1) ༩ gང།) ། ། ། ན 2 (s)) 2 ༡ (༢-) + །) (s - ད༡ -༥(5) - 6 (5) t 3 2-e at since (+3) = 1 and 2 (= Now, useing Inverse taplace transformasion both side of S-2 S (S-1) & Y S it